Ram borrowed Rs.1.50,000 from sita at the rat of 21% per year at the end of 9 months how much compound interest should be pay in nine months compounded half year.
Answer by Theo(13342)

i'm not sure what your denominations of money are.
however, assuming you are talking about rs.x, then you would do the following.

the interest rate is 21% a year compounded semi-annually.
that makes the interest rate 21/2 = 10.5% per half year = .105 per half year.

the formula is f = p * (1 + r) ^ n
f is the future value
p is the present value
r is the interest rate per time period
n = the number of time periods.

6 months is equal to one semi-annual period, therefore 9 months is equal to 1.5 semi-annual periods.

the formula of f = p * (1 + r) ^ n becomes:
f = x * (1 + .105) ^ 1.5
this becomes f = x * 1.105 ^ 1.5 which becomes:
f = x * 1.161564731.

replace x with whatever value you are talking about and that's your answer.
for example, if x = rs 1.5, then f = 1.5 * 1.161564731 = rs 1.742347097.

any questions, email to dtheophilis@gmail.com